# Sort-free Gaussian Splatting via Weighted Sum Rendering

### v1

@@ 2 Related Work
Order Dependent Transparency 3DGS The novel 3D Gaussian Splatting (3DGS) and its numerous follow-up works primarily utilize anisotropic 3D Gaussians for scene representation and an efficient differentiable rasterizer to project these Gaussians onto the image plane, either in pixel or feature representation Hamdi et al. (2024); Lin et al. (2024); Yu et al. (2024); Liang et al. (2024); Yan et al. (2024); Lu et al. (2024); Bulò et al. (2024); Fan et al. (2024); Fu et al. (2024); Zhang et al. (2024); Zou et al. (2024); Jiang et al. (2023); Qin et al. (2024); Li et al. (2024). These methods enable fast, high-resolution rendering while maintaining excellent quality. For a comprehensive overview, please refer to recent survey papers such as Chen & Wang (2024a). A crucial requirement in these methods for proper blending and rendering is the order of the Gaussians, which are typically sorted by their depth using a tile-based sorting algorithm. However, this sorting requirement introduces several challenges in practical implementation and visual quality, such as sudden changes in the appearance of object parts or “popping” artifacts, as recently addressed in Radl et al. (2024). While their results show a limited increase in computational complexity, the necessary sorting modifications are significantly more involved.
@@
where $d_{i}$ is the distance to camera, $\mathbf{c}_{0}$ is the background color, $w\left(\cdot\right)$ is a function that decreases with distance, so that objects nearer to the camera are assigned larger weights. In the OIT rendering Eq. 2 there are two sums defining pixel values, and since addition is commutative, they can be computed in any order. The final pixel value can only be computed after the two sums are known, which can be implemented as a single pass post-process over the image.
@@ 3 Preliminaries: Gaussian Splatting
Critically, due to it’s non-commutative blending operations requiring a strict depth ordering, it is necessary to sort Gaussians according to their distance to the camera position $\mathbf{f}$. Formally, it is necessary to determine a sorting sequence $s[k]$ of indexes according to increasing depth such that
$k\geq l\Rightarrow d_{s\left[k\right]}\geq d_{s\left[l\right]},k,l=1,2,\cdots,\mathcal{N}$ (6)
@@ 4.2 Weighted Sum Rendering
However, DIR-WSR does not work well for the complex scenes, and often introduces blurring artifacts in areas where Gaussians overlap. We attribute the poor visual performance due to the lack of depth information, which prevents accurate estimation of transparency. A visual example is provided in Figure 2, illustrating two non-transparent Gaussians with $\alpha=1$ with different color [^1] Please note that in WSR, $\alpha$ can be larger than 1. We assume $\alpha=1$ as non-transparency for illustrative purposes.. Since DIR-WSR uses a constant weight of 1, it is unable to correctly handle this situation, leading to an incorrect estimation of the resulting color as purple.
@@ 5 Experimental Results
Table 2: Runtime (ms) comparion using a Snapdragon 8 Gen 3 GPU on the Mip-NeRF360, Tanks & Temples and Deep Blending datasets. The images are rendered at a resolution of $1920\times 1080$. When the mobile resources are exhausted, it would significantly slow down the rasterization (labeled with “*”).
@@ 5.1 Comparison with state-of-the-art methods
We compare our method with state-of-the-art techniques, including Plenoxels Fridovich-Keil et al. (2022), INGP Müller et al. (2022), M-NeRF360 Barron et al. (2022), and 3DGS Kerbl et al. (2023). The results in Table 1 demonstrate that our method, while avoiding any sorting, achieves comparable results to the baseline 3DGS. In terms of PSNR, our method outperforms 3DGS by 0.47dB and 0.22dB on the Tanks & Temples and Deep Blending datasets, respectively, while averaging only 0.02dB less on the Mip-NeRF360 dataset. For visual comparison, Figure 4 shows that our LC-WSR excellently recovers challenging details of the fireplace in the “Dr Johnson” scenes. In other scenes, WSR particularly recovers better fidelity in areas with strong illumination and reflection.
@@ 5.2 Computational complexity study
Table 2 reports the runtime comparison of 3DGS-Compute, 3DGS-Graphics, and our LC-WSR on the Mip-NeRF360, Tanks&Temples, and Deep Blending datasets. Compared to 3DGS-Compute, 3DGS-Graphics is more efficient as it eliminates the replication steps and thus reducing the number of Gaussians for sorting. Besides 3DGS-Graphics can also better utilize the graphic fixed-functions in the hardware-supported graphic rasterization pipeline. Our method is in most cases faster than the 3DGS-Graphics method due to the elimination of the sorting step and fewer Gaussians. Note that for certain scenes using the 3DGS-Graphics method, we can exhaust our edge device’s resources due to the sorting step and high Gaussian count, which will drastically slow down the rasterization step (labeled with “*”). In contrast, our method is fully capable of running all scenes in real-time. In theory, our method’s lighting pass should be slightly slower because opacity is calculated using SH, and our rasterization step should be slower due to the weight calculations in the fragment shader and the extra subpass for weight normalization. However, we observed that WSR tends to generate models with fewer Gaussians (scene average 2.88M) compared to the original 3DGS (scene average 3.98M), which is consistent with prior work  Radl et al. (2024) that suggests modifying opacity leads to a reduction in the number of Gaussians. This also contributes to faster lighting and rasterization steps for us compared with 3DGS-Graphics. For certain scenes (kitchen and train) that have a similar number of Gaussians, our total time is close to or slightly worse compared to the original model because the overhead added in the rasterization step is not fully compensated by the removal of the sorting step, which may indicate further optimizations are needed for our method’s rasterization implementation. Overall, the proposed method is on average 1.23× faster in total time when the mobile resources are not exhausted.
Table 3 compares the memory of our method with 3DGS-Graphics and 3DGS-Compute. Compared to 3DGS-Compute, 3DGS-Graphics utilizes less memory as it performs global sorting thereby eliminating the replication step. Our method only requires around 63% memory of 3DGS-Graphics. This memory reduction can be attributed to removing the sorting operations as well as fewer total Gaussians. Specifically, removing sorting reduces around 17% memory, while the reducing the number of Gaussians contribute the remaining.
Table 3: Memory (MB) comparison on the Mip-NeRF360, Tanks & Temples and Deep Blending datasets. The images are rendered at a resolution of $1920\times 1080$.
@@ 5.3 Ablation study
Table 4: Comparison between WSR methods on the Mip-NeRF360, Tanks & Temples and Deep Blending datasets.
@@ 6 Conclusions
Limitations and Future Work. Early-Termination. In 3DGS a pixel can be easily terminated by monitoring if its accumulated alpha has been saturated. Alternately, we could implement heuristic to halt further evaluation of Gaussians whose opacity falls below some threshold, though our current WSR implementation does not support such an optimization. We implemented our method using the traditional graphics pipeline to make efficient use of mobile GPUs, especially the hardware rasterizer. In our implementation, opacity and weight computations are performed in the fragment shader, while blending is handled by the hardware, since every fragment operates independently and does not have access to the accumulated opacity, we are unable to easily implement early termination. This would require reading the render target’s content within the invoking draw pass, which necessitates costly read-modify-write operations and requires Vulkan extensions that many mobile devices do not currently support. Additionally, since the screen-space location of each Gaussian is determined only after the vertex shader, a pixel can only be discarded very late in the pipeline. Compact Gaussians. Recently, several compact Gaussian methods have been introduced to significantly reduce the number of Gaussians while maintaining visual performance Jo et al. (2024); Lee et al. (2024); Fan et al. (2023); Niedermayr et al. (2024); Chen et al. (2024); Morgenstern et al. (2023); Navaneet et al. (2023). We believe that our sort-free method can benefit from the rapid pace of research on compact Gaussians.

### v2

@@ 2 Related Work
Order Dependent Transparency 3DGS The novel 3D Gaussian Splatting (3DGS) and its numerous follow-up works primarily utilize anisotropic 3D Gaussians for scene representation and an efficient differentiable rasterizer to project these Gaussians onto the image plane, either in pixel or feature representation Hamdi et al. (2024); Lin et al. (2024); Yu et al. (2024); Liang et al. (2024); Yan et al. (2024); Lu et al. (2024); Bulò et al. (2024); Fan et al. (2024); Fu et al. (2024); Zhang et al. (2024); Zou et al. (2024); Jiang et al. (2023); Qin et al. (2024); Li et al. (2024); Jo et al. (2024); Lee et al. (2024); Fan et al. (2023); Niedermayr et al. (2024); Chen et al. (2024); Morgenstern et al. (2023); Navaneet et al. (2023). These methods enable fast, high-resolution rendering while maintaining excellent quality. For a comprehensive overview, please refer to recent survey papers such as Chen & Wang (2024a). A crucial requirement in these methods for proper blending and rendering is the order of the Gaussians, which are typically sorted by their depth using a tile-based sorting algorithm. However, this sorting requirement introduces several challenges in practical implementation and visual quality, such as sudden changes in the appearance of object parts or “popping” artifacts, as recently addressed in Radl et al. (2024). While their results show a limited increase in computational complexity, the necessary sorting modifications are significantly more involved.
@@
where $d_{i}$ is the distance to camera, $\mathbf{c}_{0}$ is the background color, $w\left(\cdot\right)$ is a function that decreases with distance, so that objects nearer to the camera are assigned larger weights. In the OIT rendering Eq. 2 there are two sums defining pixel values, and since addition is commutative, they can be computed in any order. Inspired by OIT, our method extends to Gaussian Splatting by introducing the learnable parameters and view-dependent opacity. Besides, compared to the OIT exponential weights, linear weights produce better PSNR results for Gaussian Splatting.
@@ 4.2 Weighted Sum Rendering
However, DIR-WSR does not work well for complex scenes, and often introduces blurring artifacts in areas where Gaussians overlap. We attribute the poor visual performance due to the lack of depth information, which prevents accurate estimation of transparency. A visual example is provided in Figure 2, illustrating two non-transparent Gaussians with $\alpha=1$ with different color [^1] Please note that in WSR, $\alpha$ can be larger than 1. We assume $\alpha=1$ as non-transparency for illustrative purposes.. Since DIR-WSR uses a constant weight of 1, it is unable to correctly handle this situation, leading to an incorrect estimation of the resulting color as purple.
@@ 5 Experimental Results
Table 2: Runtime (ms) comparion using a Snapdragon 8 Gen 3 GPU on the Mip-NeRF360, Tanks & Temples and Deep Blending datasets. The images are rendered at a resolution of $1920\times 1080$. (“*” indicates that mobile resources are exhausted).
@@ 5.1 Comparison with state-of-the-art methods
We compare our method with state-of-the-art techniques, including Plenoxels Fridovich-Keil et al. (2022), INGP Müller et al. (2022), M-NeRF360 Barron et al. (2022), 3DGS Kerbl et al. (2023), Compact 3DGS Morgenstern et al. (2023), Compressed 3DGS Niedermayr et al. (2024), and LightGaussian Fan et al. (2023). The results in Table 1 demonstrate that our method, while avoiding any sorting, achieves comparable results to the baseline 3DGS. In terms of PSNR, our method outperforms 3DGS by 0.47dB and 0.22dB on the Tanks & Temples and Deep Blending datasets, respectively, while averaging only 0.02dB less on the Mip-NeRF360 dataset. For visual comparison, Figure 4 shows that our LC-WSR excellently recovers challenging details of the fireplace in the “Dr Johnson” scenes. In other scenes, WSR particularly recovers better fidelity in areas with strong illumination and reflection.
@@ 5.2 Computational complexity study
Table 2 reports the runtime comparison of 3DGS-Compute, 3DGS-Graphics, and our LC-WSR on the Mip-NeRF360, Tanks&Temples, and Deep Blending datasets. Compared to 3DGS-Compute, 3DGS-Graphics is more efficient as it eliminates the replication steps and thus reducing the number of Gaussians for sorting. Besides 3DGS-Graphics can also better utilize the graphic fixed-functions in the hardware-supported graphic rasterization pipeline. Our method is in most cases faster than the 3DGS-Graphics method due to the elimination of the sorting step and fewer Gaussians. Note that for certain scenes using the 3DGS-Graphics method, we can exhaust our edge device’s resources due to the sorting step and high Gaussian count, which will drastically slow down the rasterization step (labeled with “*”). Unfortunately, it can be difficult to pinpoint the sources of inefficiency in complex GPUs. To the best of our knowledge, this is caused by sudden changes to many more cache misses. In contrast, our method is fully capable of running all scenes in real time. In theory, our method’s lighting pass should be slightly slower because opacity is calculated using SH, and our rasterization step should be slower due to the weight calculations in the fragment shader and the extra subpass for weight normalization. However, we observed that WSR tends to generate models with fewer Gaussians (scene average 2.88M) compared to the original 3DGS (scene average 3.98M), which is consistent with prior work  Radl et al. (2024) that suggests modifying opacity leads to a reduction in the number of Gaussians. This also contributes to faster lighting and rasterization steps for us compared with 3DGS-Graphics. For certain scenes (kitchen and train) that have a similar number of Gaussians, our total time is close to or slightly worse compared to the original model because the overhead added in the rasterization step is not fully compensated by the removal of the sorting step, which may indicate further optimizations are needed for our method’s rasterization implementation. Overall, the proposed method is on average 1.23× faster in total time when the mobile resources are not exhausted.
[…]
Table 3: Runtime memory (MB) comparison on the Mip-NeRF360, Tanks & Temples and Deep Blending datasets. The images are rendered at a resolution of $1920\times 1080$.
@@ 5.3 Ablation study
Table 4: WSR variants on the Mip-NeRF360, Tanks & Temples and Deep Blending datasets.
@@ Appendix B Per Scenes Evaluation Metric
Figure 6: Moving camera along $z$ direction. EXP-WSR and LC-WSR produces plausible results.
Appendix C More Discussions
Compared with gsplat webgl gsplat.tech At the same rendering resolution, our method improve the rendering speed from 19-25 fps to 30fps on a Snapdragon 8 gen3. The gsplat implementation runs at a significantly lower resolution and leverages an asynchronous CPU sort running at a frequency below the framerate, which exacerbates temporally artifacts under motion. Our implementation does not have these artifacts, which demonstrates the effectiveness of our method.
Moving camera along $z$ direction. The alpha and weights are applied equally to the RGB color components. Ratios of different splats are maintained along the z-direction with Exponential Weighted Sum Rendering (EXP-WSR), but not with Linear Correction Weighted Sum Rendering (LC-WSR). However, LC-WSR also produces visually plausible results. We conducted the experiments to move the camera along $z$ direction. In the experiments, we test our EXP-WSR and LC-WSR on the Bicycle and Garden scenes. As shown in Figure 6, we don’t observe the color shifting artifacts.
Figure 7: Failure case. The seat seems to be transparent, as it is in dark and the bicycle frame is in bright white.
Densification and pruning for view dependent opacity. The 3DGS densification technique was maintained for our method without any modification. And we removed the pruning based on the opacity threshold. There are researches exploring other signals, such as mask Lee et al. (2024), color Bulò et al. (2024), for densification and pruning, and achieving promising results. Our method will benefit from these works in the future.
Culling in WSR. Culling operates the same as 3DGS. As with 3DGS, we render only those Gaussians within the Camera Frustum. Our implementation does sum over all visible Gaussians, we configure the hardware to perform this summation efficiently using hardware supported blending operations.
Failure case. One type of failure that is easy to identify visually is apparent transparency of dark objects when in front of a very light background. Figure 7 shows an example where our method fails. However, this could be detected with differentiable rendering and thus can be attributed to suboptimal training.
Appendix D Limitations and Future Work
Early-Termination. In 3DGS a pixel can be easily terminated by monitoring if its accumulated alpha has been saturated. Alternately, we could implement heuristic to halt further evaluation of Gaussians whose opacity falls below some threshold, though our current WSR implementation does not support such an optimization. We implemented our method using the traditional graphics pipeline to make efficient use of mobile GPUs, especially the hardware rasterizer. In our implementation, opacity and weight computations are performed in the fragment shader, while blending is handled by the hardware, since every fragment operates independently and does not have access to the accumulated opacity, we are unable to easily implement early termination. This would require reading the render target’s content within the invoking draw pass, which necessitates costly read-modify-write operations and requires Vulkan extensions that many mobile devices do not currently support. Additionally, since the screen-space location of each Gaussian is determined only after the vertex shader, a pixel can only be discarded very late in the pipeline.
Compact Gaussians. Recently, several compact Gaussian methods have been introduced to significantly reduce the number of Gaussians while maintaining visual performance Jo et al. (2024); Lee et al. (2024); Fan et al. (2023); Niedermayr et al. (2024); Chen et al. (2024); Morgenstern et al. (2023); Navaneet et al. (2023). We believe that our sort-free method can benefit from the rapid pace of research on compact Gaussians.